# Assurance of AI Systems From a Dependability Perspective

### v1

@@ 3 Assurance of AI Systems for Specific Functions
In this section we consider novel non-CPS systems and applications that are made possible or are performed in new ways by the capabilities of AI and ML. These include systems that play games of skill or strategy, those that design things or perform scientific predictions such as protein folding or weather forecasting, the generation or management of responses based on these, decision support systems that analyze medical images or loan, job, and college applications or prisoner sentencing and parole and so on, and also systems that use Large Language Models (LLMs) such as ChatGPT and other generic AI and ML capabilities to perform specific functions, which can include generation of images and video as well as text. We do not include generic capabilities themselves (those are in the next section).
The reason we focus here on systems that perform specific functions is that assurance needs to identify hazards; the hazards of systems that perform specific functions can be conjectured from the functions concerned and the environment in which they are deployed, whereas the hazards of generic systems such as LLMs have to consider all the functions they might be called to perform and all the environments they might be deployed in.
@@
More likely, however, assurance must depend on some form of runtime verification, implemented either by additional training and careful prompting for “fine tuning” the LLM,181818 LLMs are adapted to specific tasks by instructions, referred to as “prompts,” given in natural language [123]. or by explicit guards. As with the autonomous systems of the previous section, runtime verification poses difficulties due to lack of non-AI ways to perceive the context or world model within which to make the guarding decisions. For example, if an LLM is suspected of racial bias despite race being absent from its training and operational data, a guard also lacking true data on race might need to have its own ML component to “perceive” this attribute.
@@
There are many proposals for constructing “guardrails” within or around LLMs. We discuss some of these in the next section on assurance for generic AI but here we will consider some that can be customized to specific applications.
One approach is to develop “workflows” (i.e., sequences of prompts) where the LLM is iteratively asked to critique and improve its previous outputs. A recent paper reports precision around 90% for a workflow that extracts data (as Material, Value, Unit triplets) from materials science papers [127] and similar approaches have been successful in other applications [104]. Another direction exploits the large “context window” (i.e., input) allowed by some recent LLMs to provide a prompt with hundreds of training examples prior to the real query [2] (the paper also proposes ways to automatically generate suitable examples). Previously, such “fine tuning” required access to the LLM’s training environment and adjustment to the weights in its neural net. Related to this are applications that provide a substantial input and then ask the LLM to do something with it (e.g., summarize it, or identify its topic).
All these approaches exploit the natural language capability of the LLM, but constrain it to operate within the input provided, so there is little opportunity to generate content extracted from its training corpus. Due to its model-free nature, the LLM may still misinterpret the input and do a wrong thing, but this should also be minimized as a large input can explicitly (via extensive prompts) or implicitly (via a block of text) convey the intended context or model.
Another approach operates by attempting to access the model or “reasoning” purportedly employed. This is Explainable AI, which aims to deliver reasons for accepting the output from an LLM. Counterfactual explanations, where a system may deliver a response such as “I am declining your loan application but would have approved it if your income was \$5,000 greater, or your deposit was \$2,000 more” [169, 45], may be particularly useful for specific applications as they should be expressed in terms of the policy specified in the prompt and not arbitrary “facts” from the training corpus. A runtime checker could reject decisions whose counterfactual explanations violate the policy. Do note, however, that counterfactual explanations can be manipulated [155], so this approach needs to be employed with caution. However, in some applications it does not matter how a decision and its explanation are produced: as long as they pass an independent check (that does incorporate models of the appropriate policy), they may be considered valid.
@@
A particular attraction of such dual process architectures is that the second (upper) process does have a (generic) model and compensates for the model-free behavior of the first. A more sophisticated dual-process architecture, as sketched in the previous section and inspired by that of the human brain, uses predictive processing. Here, the two processes combine to build an explicit model of the world at runtime [^21] We should explain how a model-free perception system is used to build a world model. The object detector of a self-driving car, say, has no model of what it is seeing: it does not “know” that it is on a road and looking at other traffic; it has simply been trained to recognize objects (albeit using billions of road images). Hence it may report a “boat” when “trailer” (with boat as cargo) would be more useful. The upper levels of the perception pipeline are explicitly designed to build a model of traffic on a road, and apply this built-in interpretation to the model-free output of the object detector. and, rather then rely directly on the model-free products of the lower process, this highly specific model is used to predict them and they are therefore interpreted in this light; the higher-level process intervenes when a large prediction error indicates “surprise.”
@@
4 Assurance for Generic AI
By generic AI we mean those AI and ML systems that are broadly capable and can be adapted to some specific application with relatively little effort. Canonically, these are generative “foundation models” such as LLMs (for language) or “generative adversarial networks” (GANs) and “diffusion models” (for images) that are pre-trained using unsupervised learning on vast amounts of unlabeled data so that they learn general features and associations rather than specific skills. They can be adapted to specific tasks using “prompts,” and “prompt engineering” has become a recognized activity [123]. We refer to such systems generically as LLMs and we discussed assurance for applications built on them in the previous section; here, we focus on the LLMs themselves.
@@ 4 Assurance for Generic AI
One manifestation of this generality and difficulty is that public discourse and much technical literature speaks of trust rather than assurance. Indeed, many current AI systems are sufficiently good at natural language, including emotional and cultural nuances, that people often anthropomorphize them [^23] For example, in June of 2022, Blake Lemoine, a Google engineer working with their large language model, LaMDA (Language Model for Dialogue Applications), claimed it had become “sentient” (Washington Post, 11 June 2022). and spontaneously endow them with trust. In fact, the developers of generic LLMs engineer them to encourage, and to some extent earn, this trust, as we now outline.
A bare LLM, trained to predict the next or missing “token” in a string of text, is as likely to produce wrong or noxious results as good ones. This is because its training uses a massive and largely indiscriminate sampling of material found on the Web, much of which is false or noxious. The bare LLM is then “protected” (from generating noxious or potentially harmful output) using reinforcement learning where human reviewers provide feedback on candidate results [11]. This technique has been developed into a more automatic process by first providing the LLM with a list of principles, referred to as its “constitution” and then using reinforcement learning with the feedback on alignment between its responses and its constitution provided by a second AI agent (a process referred to as “Reinforcement Learning with AI Feedback”, RLAIF) [10]. The constitution used with Anthropic’s LLM Claude is available online https://www.anthropic.com/news/claudes-constitution. It consists of a few dozen prompts, including the following examples.
@@
As we discussed in the previous section, most generic ML operates purely by statistical associations: it is model-free and has no understanding of the world, nor of right and wrong, true and false. Thus, in our opinion, interpersonal trust is inappropriate for AI and exposes its users to risk. The only trust that should be applied to an AI system is the social trust earned by a well-engineered technology. To explore this, we need to probe further into the concept of “social trust.”
@@
We will interpret the rules and constraints underlying social trust as contracts and thereby equate social trust with contractual trust. If we further interpret failure to uphold a contract as a hazard, and “vulnerability” as indicating that such failures impose costs on the trustor, then contractual trustworthiness looks very much like dependability assurance with the contract as its requirements. For this reason, we regard contractual trust in AI as an issue of assurance. However, contractual trust requires a contract and these are more readily understood with the focused AI systems of Sections 2 and 3 (where the contract corresponds to their dependability requirements) than with the generic systems considered here.
@@
Even if an agreed interpretation and weighting of the basic principles were built in to generic AI systems, it may not be obvious how to apply them. For example, a self-driving car might be confronted by a vehicle crossing against the lights and the choices are to crash into it, likely killing or injuring the occupants of both vehicles, or to swerve onto the sidewalk, likely killing pedestrians. The fairness principle might argue that all lives are equal and utilitarianism might then suggest a decision that minimizes the probable injuries. On the other hand, the care principle might argue that the system has a special responsibility for its own passengers and should seek a solution that minimizes their harm.
@@
Ethical and related principles are referred to as normative, and AI systems have been developed that can represent such principles and thereby perform “normative reasoning” [39]. These have been applied to trolley problems, including some that involve self-harm (e.g., throwing yourself in front of the trolley) and thereby violate the “unintended” aspect of DDE [31, 63]. It is claimed that fairly sophisticated logical treatments (e.g., intensional logics, counterfactuals, deontic modalities) are needed to represent normative scenarios, and these might be additional to what is needed for the primary functions of the system (hence, must be introduced explicitly, which will add complexity). Additionally, when normative requirements are introduced as rules in some formal notation, there is concern that they may have internal conflicts or otherwise lack wellformedness, and methods have been proposed for checking this [56].
@@
Building an accurate and comprehensive model of the world is difficult in a focused application; it is next to impossible for a generic tool lacking information on the possible contexts of its use.
@@
We conclude that general constraints built into guards for generic AI tools cannot provide automated protection nor assurance, largely because their correct application depends on the guard building an accurate model of the world in which to interpret them, or trusting the LLM itself to provide a suitable model (which is contrary to its generally model-free operation). It is, however, possible that, with suitable programming interfaces or suitable sensitivity to prompts, the presence of general constraints could provide useful capabilities that humans can use to add controls for specific applications built on these generic tools, albeit with weak assurance.
@@
Jacovi and colleagues [81] discuss “intrinsic” and “extrinsic” trust in AI systems. The latter is based on evaluation of observed behavior and we have previously, in Section 1, discussed the infeasibility of deriving high levels of assurance from observation alone. In those discussions, the obstacle was the vast number of tests needed for useful assurance in the absence of justified prior confidence. In the context of generic AI systems, less assurance may be acceptable and, given some prior confidence, the number of tests might be feasible; instead, the difficulty becomes construction and evaluation of adequately wide-ranging tests, given that potential applications are unknown. Some developers of generic tools are reported to manage this using crowdsourcing. [^28] For example, using the Amazon Mechanical Turk: https://www.mturk.com/. An alternative approach uses one AI system to test and evaluate another [64]. A reputation system, as sketched above, can be seen as a way of assigning and updating extrinsic trust in a “live” system.
@@
The developers of LLMs and other generic AI and ML systems are actively working to improve their trustworthiness [17]. We expect that we will see major and perhaps surprising developments in both capability and trustworthiness over the next few years, but we doubt this will achieve true assurance (i.e., explicit reasons for trust) and we believe that the methods presented here cover the feasible approaches for some degree of dependability assurance in applications of generic systems. Because assurance for their general behavior is difficult, we recommend that developers of LLMs and other general-purpose systems provide APIs and “hooks” so that those who construct specific applications on their foundation can program in protections or guards suitable to their context.
@@ 5 Assurance and Alignment for AGI
The public is also aware of rapid increase in the capabilities of ML for image and language recognition and in automated perception, as seen in face recognition, language translation, voice assistants, and self-driving cars. Little more than a decade ago, DARPA’s “Grand Challenge” for autonomous cars to drive a course in open country left all its participants failed or crashed [34]. Yet today, self-driving cars are routine, if not yet fully safe. The general public is less aware that AI technology also performs automated design, where ML-enabled systems rapidly generate and evaluate designs for buildings, drugs, underwater and aerial drones, and so on, and can even propose scientific theories [22] and solve open math problems [138].
@@
In previous sections we have considered assurance for specific systems and for generic tools based on AI and ML where the concern is that faulty behavior may lead to harm. For projected developments of current AI and ML systems and the potential emergence of AGI, however, the concern is not just faulty behavior as the social impact of new capabilities. In particular, an AGI system capable of setting its own goals might pursue objectives contrary to human interests. Hence, the assurance problem for AGI is to ensure that its goals align with those of human society [173]—and this needs to be maintained even though AGI may fall into the hands of bad actors, criminals, and adversaries.
@@
Some consider “safe superintelligence is the most important technical problem of our time” (https://ssi.inc), but an insidious form of disruption may arise long before AGI is available: namely, AI systems that are good enough and cheap enough to displace (possibly superior) human services—what we might call Artificial Fairly Good Intelligence (AFGI). For example, well-researched journalism already finds it hard to compete with LLM-generated press releases and propaganda that masquerade as “news.” Related to this is another near-term disruption: the possibility that ubiquitous use of LLMs and related AI tools will reduce our collective sagacity by repeatedly circulating misinformation, falsehoods, and mediocrity, so that we lose the ability to access or recognize truth, novelty, or insight [112]. This extends to images, sound, and videos, where “deepfakes” enabled by diffusion models invert the adage that “seeing (and hearing) is believing.”
@@
Furthermore, LLMs and diffusion models are trained by scraping text, images, and all manner of other material from the Web (likely violating copyright in many instances [^30] New York Times, 27 December 2023.) and as they become widely used, their own output becomes a large part of the web corpus that trains the next generation and so on, potentially causing the Web to become full of LLM-generated pablum. There already is evidence for this: replying to a prompt asking it to tinker with malware, an LLM called Grok, developed by xAI, refused “as it goes against OpenAI’s use case policy.” OpenAI has nothing to do with xAI, so Grok presumably generated this response by scraping web text generated by OpenAI’s LLMs. This phenomenon is termed “Model Collapse” [153] and is predicted to “dumb down” the overall information content of the Web. Meanwhile, recommender systems for music, video, and books may lead us to serendipitous and enjoyable discoveries, but they can also be manipulated to channel our attention along predetermined paths.
@@
A generally anticipated disruption is widespread unemployment when capabilities a little better than those available today, but still far short of AGI, displace office jobs ranging from clerks to middle-management, software coding, routine design, and professions such as lawyer and doctor. These AFGI capabilities might be superior to human performance, or inferior but still “good enough,” and cheaper; either way, they could displace human jobs. Beyond harm to individuals, unfettered use of AFGI, particularly in conjunction with polarization perpetrated by social media, could undermine our institutions, or our trust in these. In the hands of adversaries, AFGI-enabled disinformation campaigns and social media manipulations could sow societal discord and conflict. Until recently, interventions of this kind required the capabilities and resources of a nation state, but are now coming within reach of lone actors. All these disruptions could be chronic rather than acute: that is, they could develop over a period (e.g., creeping unemployment) rather than abruptly and, as with climate change, this may make recognition difficult, and effective response hard to mobilize.
@@
In addition to these cases, where harm may be unintended, there are dystopian applications of AFGI such as “Lethal Autonomous Weapons Systems” (LAWS, aka. “killer robots”). Most countries have policies that require some form of human control or approval over these (e.g., USA [152]) but it is unknown how these limitations will hold up in combat. And there are AI systems that contribute to lethality without being weapons. For example, AI-powered intelligence systems identify vastly more potential targets more quickly than humans can (e.g., 100 a day vs. 50 a year [9]). Humans are required to select and approve actual targets from those identified but the sheer numbers make meaningful oversight difficult.
@@
These steps are well-intentioned but apart from the EU’s GPAI proposals, we consider them unlikely to provide much benefit. Every societal hazard posed by recent computer systems was unanticipated and unrecognized until the harm was done, and subsequent regulation has proved ineffective or counterproductive. For example, the advent of advertising as the main source of revenue for Web services has led to universal surveillance of personal online activity, to the extent that even our cars now invade our privacy. [^31] See https://foundation.mozilla.org/en/privacynotincluded/categories/cars/. This was not anticipated and regulatory responses such as “cookie warnings” are a source of annoyance rather than protection, and indicate technological illiteracy on the part of regulatory agencies and lawmakers. Similarly, the poisonous impact of “algorithmic” traffic generation on social media was not anticipated and still has no effective response.

### v2

@@ 3 Assurance of AI Systems for Specific Functions
In this section we consider novel non-CPS systems and applications that are made possible or are performed in new ways by the capabilities of AI and ML. These include systems that play games of skill or strategy, those that design things or perform scientific predictions such as protein folding or weather forecasting, the generation or management of responses based on these, decision support systems that analyze medical images or loan, job, and college applications or prisoner sentencing and parole and so on, and also systems that use Large Language Models (LLMs) such as ChatGPT and other general-purpose AI and ML capabilities to perform specific functions, which can include generation of images and video as well as text. We do not include general-purpose capabilities themselves (those are in the next section).
The reason we focus here on systems that perform specific functions is that assurance needs to identify hazards; the hazards of systems that perform specific functions can be conjectured from the functions concerned and the environment in which they are deployed, whereas the hazards of general-purpose systems such as LLMs have to consider all the functions they might be called to perform and all the environments they might be deployed in.
@@
More likely, however, assurance must depend on some form of runtime verification, implemented either by additional training and careful prompting for “fine tuning” the LLM,191919 LLMs are adapted to specific tasks by instructions, referred to as “prompts,” given in natural language [127]. or by explicit guards. As with the autonomous systems of the previous section, runtime verification poses difficulties due to lack of non-AI ways to perceive the context or world model within which to make the guarding decisions. For example, if an LLM is suspected of racial bias despite race being absent from its training and operational data, a guard also lacking true data on race might need to have its own ML component to “perceive” this attribute.
@@
There are many proposals for constructing “guardrails” within or around LLMs. We discuss some of these in the next section on assurance for general-purpose AI but here we will consider some that can be customized to specific applications.
One approach is to develop “workflows” (i.e., sequences of prompts) where the LLM is iteratively asked to critique and improve its previous outputs. A recent paper reports precision around 90% for a workflow that extracts data (as Material, Value, Unit triplets) from materials science papers [131] and similar approaches have been successful in other applications [108]. Another direction exploits the large “context window” (i.e., input) allowed by some recent LLMs to provide a prompt with hundreds of training examples prior to the real query [2] (the paper also proposes ways to automatically generate suitable examples). This is called “in-context learning” [34]; previously, such “fine tuning” required access to the LLM’s training environment and made adjustment to the weights in its neural net. Related to this are applications that provide a substantial input and then ask the LLM to do something with it (e.g., summarize it, or identify its topic).
All these approaches exploit the natural language capability of the LLM, but constrain it to operate on or within the input provided, so there is little opportunity to generate or insert content extracted from its training corpus. Due to its model-free nature, the LLM may still misinterpret the input and do a wrong thing, but this should also be minimized as a large input can explicitly (via extensive prompts) or implicitly (via a block of text) convey the intended context or model.
Another approach operates by attempting to access the model or “reasoning” purportedly employed. This is Explainable AI, which aims to deliver reasons for accepting the output from an LLM. Counterfactual explanations, where a system may deliver a response such as “I am declining your loan application but would have approved it if your income was \$5,000 greater, or your deposit was \$2,000 more” [173, 47], may be particularly useful for specific applications as they should be expressed in terms of the policy specified in the prompt and not arbitrary “facts” from the training corpus. A runtime checker could reject decisions whose counterfactual explanations violate the policy. Do note, however, that counterfactual explanations can be manipulated [159], so this approach needs to be employed with caution. However, in some applications it does not matter how a decision and its explanation are produced: they can be considered valid as long as they pass an independent check (that does incorporate models of the appropriate policy).
@@
A particular attraction of such dual process architectures is that the second (upper) process does have a (generic) model and compensates for the model-free behavior of the first. A more sophisticated dual-process architecture, as sketched in the previous section and inspired by that of the human brain, uses predictive processing. Here, the two processes combine to build an explicit model of the world at runtime [^22] We should explain how a model-free perception system is used to build a world model. The object detector of a self-driving car, say, has no model of what it is seeing: it does not “know” that it is on a road and looking at other traffic; it has simply been trained to recognize objects (albeit using billions of road images). Hence it may report a “boat” when “trailer” (with a boat as cargo) would be more useful. The upper levels of the perception pipeline are explicitly designed to build a model of traffic on a road, and apply this built-in interpretation to the model-free output of the object detector. and, rather then rely directly on the model-free products of the lower process, this highly specific model is used to predict them and they are therefore interpreted in this light; the higher-level process intervenes when a large prediction error indicates “surprise.”
@@
A legitimate question is how can we build indefeasible assurance cases for systems with AI and ML where the mechanisms and evidence for assurance are relatively weak? First, it is important to remember that indefeasible does not mean inerrant (i.e., “no failures”), it means no new information would change our judgement, and this in turn means that we have diligently searched for all hazards and all defeaters to the assurance case, and have dealt with them. But in systems with AI and ML, the total space of behaviors is unknown and so the true failure rate cannot be determined, The best that we can do is to show that the failure rate is above some conservative “floor” and that if experience reveals previously unanticipated misbehavior, then we will detect and correct it before it causes unacceptable harm.
Hence, the top claim of a system with AI and ML elements will not be “no catastrophic failure in entire deployment lifetime” (as it is for aircraft software) but something like “incident rate caused by the system shall be no more than $10^{-n}$ per hour, and potential violations experienced in operation will be detected and mitigated before a second incident.”
Such a claim will be established by evidence that includes analysis and experimentally-based reliability estimation to justify the modest “floor” plus robust fleet-wide methods to detect and record incipient failures during operation (e.g., following a “surprising” prediction error or, in the worst case, an accident) and to rapidly develop and distribute repairs (recall the discussion in Section 1.2 of repairs following incidents and precursors). Mechanisms for runtime verification such as guards are the natural location for assessing dependability weakness and detecting incipient failures, which Johansson and Koopman refer to as Safety Performance Indicators (SPI) [86]. Notice that these guards are not only monitoring the system, but also collecting evidence that will update the assurance case; we refer to this architecture as providing a dynamic assurance case. Note that it may be reasonable to use strict dependability assurance for some properties (e.g., forward collisions) and dynamic assurance for others (e.g., sideways collisions).
4 Assurance for General-Purpose AI
By general-purpose AI we mean those AI and ML systems that are broadly capable and can be adapted to some specific application with relatively little effort. Canonically, these are generative “foundation models” such as LLMs (for language) or “generative adversarial networks” (GANs) and “diffusion models” (for images) that are pre-trained using unsupervised learning on vast amounts of unlabeled data so that they learn general features and associations rather than specific skills. They can be adapted to specific tasks using “prompts,” and “prompt engineering” has become a recognized activity [127]. We refer to such systems generically as LLMs and we discussed assurance for applications built on them in the previous section; here, we focus on the LLMs themselves.
@@ 4 Assurance for General-Purpose AI
One manifestation of this generality and difficulty is that public discourse and much technical literature speaks of trust rather than assurance. Indeed, many current AI systems are sufficiently good at natural language, including emotional and cultural nuances, that people often anthropomorphize them [^24] For example, in June of 2022, Blake Lemoine, a Google engineer working with their large language model, LaMDA (Language Model for Dialogue Applications), claimed it had become “sentient” (Washington Post, 11 June 2022). and spontaneously endow them with trust. In fact, the developers of general-purpose LLMs engineer them to encourage, and to some extent earn, this trust, as we now outline.
A bare LLM, trained to predict the next or missing “token” in a string of text, is as likely to produce wrong or noxious results as good ones. This is because its training uses a massive and largely indiscriminate sampling of material found on the Web, much of which is false or noxious. The bare LLM is then “protected” (from generating noxious or potentially harmful output) using reinforcement learning where human reviewers provide feedback on candidate results [11]. This technique has been developed into a more automatic process by first providing the LLM with a list of principles, referred to as its “constitution,” and then using reinforcement learning with the feedback on alignment between its responses and its constitution provided by a second AI agent (a process referred to as “Reinforcement Learning with AI Feedback”, RLAIF) [10]. The constitution used with Anthropic’s LLM Claude is available online https://www.anthropic.com/news/claudes-constitution. It consists of a few dozen prompts, including the following examples.
@@
As we discussed in the previous section, most general-purpose ML operates purely by statistical associations: it is model-free and has no understanding of the world, nor of right and wrong, true and false. Thus, in our opinion, interpersonal trust is inappropriate for AI and exposes its users to risk. The only trust that should be applied to an AI system is the social trust earned by a well-engineered technology. To explore this, we need to probe further into the concept of “social trust.”
@@
We will interpret the rules and constraints underlying social trust as contracts and thereby equate social trust with contractual trust. If we further interpret failure to uphold a contract as a hazard, and “vulnerability” as indicating that such failures impose costs on the trustor, then contractual trustworthiness looks very much like dependability assurance with the contract as its requirements. For this reason, we regard contractual trust in AI as an issue of dependability assurance. However, contractual trust requires a contract and these are more readily understood with the focused AI systems of Sections 2 and 3 (where the contract corresponds to their dependability requirements) than with the general-purpose systems considered here.
@@
Even if an agreed interpretation and weighting of the basic principles were built in to general-purpose AI systems, it may not be obvious how to apply them. For example, a self-driving car might be confronted by a vehicle crossing against the lights and the choices are to crash into it, likely killing or injuring the occupants of both vehicles, or to swerve onto the sidewalk, likely killing pedestrians. The fairness principle might argue that all lives are equal and utilitarianism might then suggest a decision that minimizes the probable injuries. On the other hand, the care principle might argue that the system has a special responsibility for its own passengers and should seek a solution that minimizes their harm. And a different interpretation of fairness might say that the pedestrians are not participating in car travel and did not sign up to its risks, so they should be spared.
@@
Ethical and related principles are referred to as normative, and symbolic AI systems have been developed that can represent such principles and thereby perform “normative reasoning” [41]. These have been applied to trolley problems, including some that involve self-harm (e.g., throwing yourself in front of the trolley) and thereby violate the “unintended” aspect of DDE [32, 65]. It is claimed that fairly sophisticated logical treatments (e.g., intensional logics, counterfactuals, deontic modalities) are needed to represent normative scenarios, and these might be additional to what is needed for the primary functions of the system (hence, must be introduced explicitly, which will add complexity). Additionally, when normative requirements are introduced as rules in some formal notation, there is concern that they may have internal conflicts or otherwise lack wellformedness, and methods have been proposed for checking this [58].
@@
Building an accurate and comprehensive model of the world is difficult in a focused application; it is next to impossible for a general-purpose tool lacking information on the possible contexts of its use.
@@
We conclude that general constraints built into guards for general-purpose AI tools cannot provide automated protection nor assurance, largely because their correct application depends on the guard building an accurate model of the world in which to interpret them, or on trusting the LLM itself to provide a suitable model (which is contrary to its generally model-free operation). It is, however, possible that, with suitable programming interfaces or suitable sensitivity to prompts, the presence of general constraints could provide useful capabilities that humans can use to add controls for specific applications built on these general-purpose tools, albeit with weak assurance.
@@
Jacovi and colleagues [83] discuss “intrinsic” and “extrinsic” trust in AI systems. The latter is based on evaluation of observed behavior and we have previously, in Section 1, discussed the infeasibility of deriving high levels of assurance from observation alone. In those discussions, the obstacle was the vast number of tests needed for useful assurance in the absence of justified prior confidence. In the context of general-purpose AI systems, less assurance may be acceptable and, given some prior confidence, the number of tests might be feasible; instead, the difficulty becomes construction and evaluation of adequately wide-ranging tests, given that potential applications are unknown. Some developers of general-purpose tools are reported to manage this using crowdsourcing. [^29] For example, using the Amazon Mechanical Turk: https://www.mturk.com/. An alternative approach uses one AI system to test and evaluate another [66]. A reputation system, as sketched above, can be seen as a way of assigning and updating extrinsic trust in a “live” system.
@@
The developers of LLMs and other general-purpose AI and ML systems are actively working to improve their trustworthiness [17]. We expect that we will see major and perhaps surprising developments in both capability and trustworthiness over the next few years, but we doubt this will achieve true assurance (i.e., explicit reasons for trust) and we believe that the methods presented here cover the feasible approaches for some degree of dependability assurance in applications of general-purpose systems. Because assurance for their general behavior is difficult, we recommend that developers of LLMs and other general-purpose systems provide APIs and “hooks” that those who construct applications on their foundation can use to program protections or guards suitable to their specific context.
@@ 5 Assurance and Alignment for AGI
The public is also aware of rapid increase in the capabilities of ML for image and language recognition and in automated perception, as seen in face recognition, language translation, voice assistants, and self-driving cars. Two decades ago, DARPA’s “Grand Challenge” for autonomous cars to drive a course in open country left all its participants failed or crashed [36]. Yet today, self-driving cars are routine, if not yet fully safe. The general public is less aware that AI technology also performs automated design, where ML-enabled systems rapidly generate and evaluate designs for buildings, drugs, underwater and aerial drones and so on, and can even propose scientific theories [23] and solve open math problems [141].
@@
In previous sections we have considered assurance for specific systems and for general-purpose tools based on AI and ML where the concern is that faulty behavior may lead to harm. For projected developments of current AI and ML systems and the potential emergence of AGI, however, the concern is not just faulty behavior but the social impact of new capabilities. In particular, an AGI system capable of setting its own goals might pursue objectives contrary to human interests. Notice that although the terminology is seldom employed in discussion of these topics, these are nonetheless dependability failures: an AI system with potentially contrary goals is a hazard that should be recognized and should be furnished with requirements to mitigate the danger, together with dependability requirements analysis to show that they do so. However, current practice in the field frames the assurance problem for AGI as ensuring that its goals align with those of human society [177]—and this needs to be maintained even though AGI may fall into the hands of bad actors, criminals, and adversaries.
@@
Some consider that “safe superintelligence is the most important technical problem of our time” (https://ssi.inc), but an insidious form of disruption may arise long before AGI is available: namely, AI systems that are good enough and cheap enough to displace (possibly superior) human services—what we might call Artificial Fairly Good Intelligence (AFGI). For example, well-researched journalism already finds it hard to compete with LLM-generated press releases and propaganda that masquerade as “news.” Related to this is another near-term disruption: the possibility that ubiquitous use of LLMs and related AI tools will reduce our collective sagacity by repeatedly circulating misinformation, falsehoods, and mediocrity, so that we lose the ability to access or recognize truth, novelty, or insight [116]. This extends to images, sound, and videos, where “deepfakes” enabled by diffusion models invert the adage that “seeing (and hearing) is believing.”
@@
Furthermore, LLMs and diffusion models are trained by scraping text, images, and all manner of other material from the Web (likely violating copyright in many instances [^31] New York Times, 27 December 2023.) and as they become widely used, their own output becomes a large part of the web corpus that trains the next generation and so on, potentially causing the Web to become full of LLM-generated pablum. There already is evidence for this: replying to a prompt asking it to tinker with malware, an LLM called Grok, developed by xAI, refused “as it goes against OpenAI’s use case policy.” OpenAI has nothing to do with xAI, so Grok presumably generated this response by scraping web text generated by OpenAI’s LLMs. This phenomenon is termed Model Collapse [157] and is predicted to “dumb down” the overall information content of the Web. Meanwhile, recommender systems for music, video, and books may lead us to serendipitous and enjoyable discoveries, but they can also be manipulated to channel our attention along predetermined paths.
@@
A generally anticipated disruption is widespread unemployment when capabilities a little better than those available today, but still far short of AGI, displace office jobs ranging from clerks to middle-management, software coding, routine design, and professions such as lawyer and doctor. These AFGI capabilities might be superior to human performance, or inferior but cheaper and still “good enough”; either way, they could displace human jobs. Beyond harm to individuals, unfettered use of AFGI, particularly in conjunction with polarization perpetrated by social media, could undermine our institutions, or our trust in these. All these disruptions could be chronic rather than acute: that is, they could develop over a period (e.g., creeping unemployment) rather than abruptly and, as with climate change, this may make recognition difficult, and effective response hard to mobilize.
@@
In addition to these cases, where harm may be unintended, there is the threat that AFGI could multiply the capabilities of adversaries and hackers (e.g., by automating the coding of viruses, or the generation of phishing emails and provocative social media posts). In the hands of adversaries, AFGI-enabled disinformation campaigns and social media manipulations could sow societal discord and conflict. Until recently, interventions of this kind required the capabilities and resources of a nation state, but are now coming within reach of lone actors. Then there are truly dystopian applications of AFGI such as “Lethal Autonomous Weapons Systems” (LAWS, aka. “killer robots”). Most countries have policies that require some form of human control or approval over these (e.g., USA [156]) but it is unknown how these limitations will hold up in combat. And there are AI systems that contribute to lethality without being weapons. For example, AI-powered intelligence systems identify vastly more potential targets more quickly than humans can (e.g., 100 a day vs. 50 a year [9]). Humans are required to select and approve actual targets from those identified but the sheer numbers make meaningful oversight difficult.
@@
These steps are well-intentioned but apart from the EU’s GPAI proposals, we consider them unlikely to provide much benefit. Every societal hazard posed by recent computer systems was unanticipated and unrecognized until the harm was done, and subsequent regulation has proved ineffective or counterproductive. For example, the advent of advertising as the main revenue source for Web services has led to universal surveillance of personal online activity, to the extent that even our cars now invade our privacy. [^32] See https://foundation.mozilla.org/en/privacynotincluded/categories/cars/. This was not anticipated and regulatory responses such as “cookie warnings” are a source of annoyance rather than protection, and indicate technological illiteracy on the part of regulatory agencies and lawmakers. Similarly, the poisonous impact of “algorithmic” traffic generation on social media was not anticipated and still has no effective response.